Find an equation of a line when it passes throught he points (4/3, -1/2) and
(-2/3, 4).

Answer by jake_6233(74)

Find an equation of a line when it passes throught he points (4/3, -1/2) and
(-2/3, 4).
An easy way to solve this problem is to find the slope y2-y1/x2-x1 These are subscripts not exponents 4- (-1/2)/(-2/3 -4/3) = (9/2)/(-2) = -9/4 =m
Then choosing one pair of coordinates solve the equation y-y1 =m(x-x1)
y - (-1/2) = -9/4(x-4/3)
y+1/2 = -9/4(x-4/3)
y= -9/4x + 36/12 -6/12
y= -9/4x -5/2 is the answer please let me know if you need any clarification